### v3

# ANSR-DT: An Adaptive Neuro-Symbolic Learning and Reasoning Framework for Digital Twins

Abstract
In this paper, we propose an Adaptive Neuro-Symbolic Learning and Reasoning Framework for digital twin technology called “ANSR-DT.” Digital twins in industrial environments often struggle with interpretability, real-time adaptation, and human input integration. Our approach addresses these challenges by combining CNN-LSTM dynamic event detection with reinforcement learning and symbolic reasoning to enable adaptive intelligence with interpretable decision processes. This integration enhances environmental understanding while promoting continuous learning, leading to more effective real-time decision-making in human-machine collaborative applications. We evaluated ANSR-DT on synthetic industrial data, observing significant improvements over traditional approaches, with up to 99.5% accuracy for dynamic pattern recognition. The framework demonstrated superior adaptability with extended reinforcement learning training, improving explained variance from 0.447 to 0.547. Future work aims at scaling to larger datasets to test rule management beyond the current 14 rules. Our open-source implementation promotes reproducibility and establishes a foundation for future research in adaptive, interpretable digital twins for industrial applications.
Impact Statement—Digital twins are increasingly deployed in industrial environments to monitor and optimize complex systems [1]. However, current implementations often operate as black boxes, limiting trust and human-machine collaboration. Our ANSR-DT framework addresses this critical gap by integrating neural networks, symbolic reasoning, and reinforcement learning to create interpretable, adaptive digital twins. The framework achieves 99.5% accuracy in dynamic event detection while maintaining human-readable decision processes through extracted symbolic rules. This approach enables real-time adaptation to changing conditions while preserving interpretability, a crucial requirement for safety-critical applications like manufacturing systems and autonomous vehicles. ANSR-DT’s open-source implementation provides a foundation for developing trustworthy AI systems that can explain their decisions to human operators, potentially reducing operational risks and improving collaborative decision-making in industrial settings.
I Introduction
Digital Twins (DTs), real-time digital replicas of physical systems, bridge the gap between physical and digital worlds to improve monitoring, decision-making, and optimization in complex scenarios such as smart factories, autonomous systems, and adaptable environments [2, 3, 4]. Despite their promise, traditional DT implementations struggle with adaptability and interpretability, fundamentally limiting trust in safety-critical industrial settings [5]. While Dynamic Digital Twins (DDTs) and Dynamic Data-Driven Application Systems (DDDAS) attempt to address these challenges through real-time data integration [6], they remain constrained by their inability to seamlessly incorporate human input and ensure transparent decision processes [7].
The core technical challenge lies in integrating real-time adaptability with interpretability—traditional models fail because they rely on static rule sets and lack continuous learning mechanisms [8]. Current approaches suffer from three critical limitations: inability to adapt to real-time human input with response latencies under 1 second, opaque decision processes that lack transparency for human operators, and static symbolic representations that cannot evolve with changing operational conditions. Existing solutions either prioritize neural learning or symbolic reasoning, but rarely integrate them effectively to achieve both high performance and interpretability [9].
To overcome these fundamental limitations, we propose ANSR-DT, an Adaptive Neuro-Symbolic Learning and Reasoning Framework for digital twin technology. ANSR-DT combines CNN-LSTM neural networks for dynamic event detection, Prolog-based symbolic reasoning for interpretable decision-making, and Proximal Policy Optimization (PPO) reinforcement learning [10] for continuous adaptation. This integration leverages attention mechanisms [11] to prioritize critical temporal patterns while maintaining logical clarity through symbolic reasoning [12], creating an adaptive system that addresses user needs without sacrificing interpretability.
The framework makes four principal contributions to adaptive digital twin technology:
  • A hybrid neuro-symbolic architecture that integrates deep learning and symbolic reasoning to improve decision-making processes in operational environments, addressing the interpretability gap in current DT implementations [13].
  • A continuous learning mechanism through PPO-enhanced CNN-LSTM integration that ensures real-time adaptation to user preferences and environmental changes [14].
  • Demonstrated performance improvements over traditional approaches, achieving 99.5% peak validation accuracy for dynamic event detection while extracting 14 stable symbolic rules that provide logically reasoned, human-interpretable explanations.
  • An open-source implementation that facilitates reproducibility and establishes a foundation for future research in adaptive, interpretable digital twins.
The remainder of this paper proceeds as follows. Section II examines related work and positions ANSR-DT within the current research landscape. Section III details the three-layer architecture comprising physical sensor integration, neuro-symbolic processing, and reinforcement learning adaptation. Section IV describes the technical realization including synthetic data generation and component integration. Section V presents comprehensive experimental evaluation demonstrating performance improvements and adaptability enhancements. Section VI analyzes the implications and limitations of our approach. Section VII concludes with future research directions.
II Background and Related Work
Traditional digital twin implementations face persistent challenges that limit their effectiveness in dynamic industrial environments. Unlike static digital models [15], contemporary DT systems require both adaptability to changing conditions and interpretability for human operators—requirements that existing approaches struggle to satisfy simultaneously. While recent work in neuro-symbolic AI [16, 17, 18] and reinforcement learning [19, 20] offers promising directions, these advances have not been systematically integrated into digital twin architectures tailored for industrial applications.
Current DT implementations suffer from fundamental architectural limitations. Most systems rely on predefined rules that cannot evolve with changing operational conditions, requiring manual intervention whenever environmental dynamics shift beyond initial specifications [21]. This rigidity becomes problematic in industrial settings where process parameters, equipment characteristics, and operational objectives continuously evolve. Furthermore, many state-of-the-art systems operate as opaque decision-makers, making their reasoning processes difficult to understand or validate [22]. This opacity directly undermines trust, particularly when human operators must verify system recommendations before implementing critical adjustments. The limited integration of human feedback exacerbates these challenges—without effective mechanisms to incorporate operator expertise and preferences, DT systems cannot align their behavior with organizational knowledge accumulated over years of operation [23].
The adaptability deficit manifests most clearly during operational deviations. Current digital twins struggle to respond quickly to dynamic changes in the physical environment, often failing to detect subtle precursors to critical failures or recognize complex multi-sensor patterns indicating emerging problems [5]. This limitation stems partly from the isolated application of techniques—approaches tend to focus exclusively on either neural learning for pattern recognition or symbolic reasoning for rule-based inference, failing to leverage the complementary strengths of both paradigms [9]. Neural networks excel at extracting complex patterns from high-dimensional sensor data but provide little insight into their decision-making processes. Conversely, symbolic systems offer transparent reasoning but lack the flexibility to adapt to novel patterns not explicitly encoded in their rule bases.
Recent neuro-symbolic AI research demonstrates the potential of hybrid architectures that combine neural learning with symbolic reasoning. Schmidt et al. [16] systematically examined neuro-symbolic approaches for knowledge graph construction in manufacturing, identifying key integration patterns between subsymbolic and symbolic representations. Zhou et al. [17] demonstrated practical neuro-symbolic AI deployment at Bosch, highlighting the importance of data foundations and systematic integration strategies. These works establish that effective neuro-symbolic integration requires careful attention to the bidirectional flow of information between neural and symbolic components—a principle central to ANSR-DT’s architecture.
ANSR-DT builds on these foundations by integrating neural networks and symbolic reasoning specifically for digital twin adaptability and interpretability. The framework extends beyond existing neuro-symbolic approaches through its incorporation of reinforcement learning for continuous policy adaptation, enabling the system to refine both its neural pattern recognition and symbolic rule base based on observed outcomes. This three-way integration—neural learning, symbolic reasoning, and reinforcement learning—distinguishes ANSR-DT from prior work while addressing the core limitations of traditional digital twin implementations. By maintaining a streamlined dataset design and extended training protocols, the framework balances computational efficiency with the depth of adaptation required for industrial deployment.
III Proposed Framework
III-A ANSR-DT Architecture Overview
ANSR-DT integrates neural networks, symbolic reasoning, and reinforcement learning through a three-layer architecture designed to provide both adaptability and interpretability in digital twin applications. As illustrated in Fig. 1, the framework comprises a Physical Layer for sensor integration, a Processing Layer for neuro-symbolic reasoning, and an Adaptation Layer for continuous learning and rule updating. This architecture facilitates seamless integration between physical systems and their digital representations while maintaining transparent decision processes that human operators can understand and validate.
Figure 1: Overview of the ANSR-DT framework architecture. The framework consists of three main layers: (1) Physical Industrial Environment for sensor integration and human operator interaction, (2) Processing Layer implementing the neuro-symbolic reasoning engine with deep learning and symbolic components, and (3) Adaptation Layer incorporating reinforcement learning and dynamic rule updating mechanisms. Solid arrows indicate primary data flow, while dashed arrows represent feedback loops and rule updates. The framework enables real-time human-machine collaboration through continuous adaptation and interpretable decision-making.
Each layer serves a distinct function while maintaining well-defined interfaces with adjacent layers. The Physical Layer collects and preprocesses sensor data, transforming raw measurements into standardized inputs for downstream processing. The Processing Layer translates these signals into meaningful patterns through neural networks and converts detected patterns into human-readable symbolic rules through logical inference. The Adaptation Layer continuously refines the system’s decision-making policies based on performance feedback and changing operational conditions, ensuring that both neural and symbolic components evolve together. The operational sequence and information flow between these layers is detailed in Fig. 2, which shows how sensor data progresses through pattern detection, symbolic reasoning, and adaptive control decisions.
III-B Physical Layer: Sensor Integration
The Physical Layer establishes the foundation for ANSR-DT by connecting the digital twin to its physical counterpart through multiple sensor modalities:
  • Temperature Sensors monitor thermal variations in industrial equipment, detecting potential overheating conditions or thermal deviations that may indicate degraded performance or impending component failures.
  • Vibration Sensors measure mechanical oscillations, providing early warnings of component wear, imbalance, or structural degradation through characteristic frequency signatures.
  • Pressure Sensors track fluid dynamics within systems, identifying potential leaks, flow restrictions, or anomalous pressure buildup that could compromise operational safety.
Raw sensor data undergoes systematic preprocessing before entering the Processing Layer. This preprocessing pipeline includes noise filtering to remove measurement artifacts, normalization to ensure consistent scaling across sensor types, and temporal alignment to synchronize multi-sensor data streams with potentially different sampling rates [24, 25]. The standardized sensor outputs then flow into the neural component for pattern extraction, establishing a consistent data foundation for downstream reasoning.
III-C Processing Layer: Neural-Symbolic Reasoning
The Processing Layer constitutes the cognitive core of ANSR-DT, combining deep learning for temporal pattern extraction with symbolic reasoning for interpretable decision-making. This dual-component architecture addresses the fundamental tension between learning capacity and explainability that limits many contemporary AI systems.
III-C1 Neural Component: CNN-LSTM Architecture
The neural component employs a CNN-LSTM architecture comprising two Conv1D layers with 64 and 128 filters (kernel size 3), followed by batch normalization and max-pooling to extract spatial features. Two LSTM layers with 100 and 50 units then capture temporal dependencies across sensor sequences, while an attention mechanism prioritizes critical time steps within the temporal patterns. This configuration totals 189,481 parameters, carefully balanced to provide sufficient model capacity for the 5,000-sample dataset without risking overfitting on the relatively compact training set.
The neural component transforms raw sensor data into feature vectors representing detected patterns, critical events, or state changes. These vectors capture complex multivariate relationships across temperature, vibration, and pressure measurements that would be difficult to encode through manual feature engineering. The learned representations then serve as inputs to the symbolic reasoning engine, establishing a bridge between subsymbolic pattern recognition and symbolic logical inference.
III-C2 Symbolic Component: Rule-Based Reasoning
The symbolic reasoning engine translates neural outputs into interpretable logical rules and facts using a Prolog-compatible representation [26]. This component maintains a knowledge base representing system states, operational constraints, and decision logic in first-order predicate logic format. The symbolic representation provides human-readable encoding of system knowledge, explicit representation of domain expertise and operational constraints, traceable decision paths from observations to actions, and direct mechanisms to incorporate human feedback and preferences into the reasoning process.
The integration between neural and symbolic components follows the rule extraction process governed by Eq. 1. New rules $R_{\text{new}}$ are selected when they meet two criteria: derivation from neural model outputs through feature extraction function $\mathcal{F}$ operating on model parameters $M_{\theta}$ and data $\mathcal{D}$, and confidence scores $\text{conf}(r)$ exceeding threshold $\tau$. This dual-threshold approach ensures that only reliable and relevant rules enter the knowledge base.
$R_{\text{new}}=\{r\mid\text{conf}(r)>\tau,r\in\mathcal{F}(M_{\theta},\mathcal{D})\}$ (1)
In the implementation, the SymbolicReasoner component performs this integration through the extract_rules_from_neural_model function, which analyzes patterns detected by the neural network and converts them into symbolic rules. The update_rules method then incorporates these rules into the Prolog rule base for logical inference [27]. This bidirectional interface ensures that insights from neural learning inform symbolic reasoning while symbolic constraints guide neural attention toward operationally relevant patterns.
III-D Adaptation Layer: Reinforcement Learning
The Adaptation Layer enables ANSR-DT to continuously improve its performance through reinforcement learning, specifically using the Proximal Policy Optimization (PPO) algorithm to refine decision-making policies while ensuring stable updates.
III-D1 PPO Algorithm Implementation
The PPO algorithm optimizes the policy for decision-making through a clipped objective function that balances learning progress with stability, as formalized in Eq. 2. The probability ratio $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ compares the current policy’s action selection probability to the previous policy’s probability for action $a_{t}$ in state $s_{t}$. The Advantage function $A_{t}=Q(s_{t},a_{t})-V(s_{t})$ quantifies the relative value of taking specific actions compared to the average value of state $s_{t}$. The entropy term $H[\pi_{\theta}]$ promotes exploration of the action space, with its impact controlled by coefficient $\beta=0.01$ to balance exploration with exploitation of known good policies.
$L^{PPO}(\theta)=\mathbb{E}_{t}\Big[\min\Big(r_{t}(\theta)A_{t},\\ \text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\Big)-\beta H[\pi_{\theta}]\Big]$ (2)
Our implementation employs $\beta=0.01$ to balance exploration with exploitation and uses a clipping range $\epsilon=0.2$ to limit the magnitude of policy updates, preventing destabilizing changes. These parameters were selected to ensure stable learning while maintaining sufficient adaptability for dynamic industrial environments. Extending PPO training from 5 iterations (10,240 timesteps) to 98 iterations (200,704 timesteps) improved explained variance from 0.447 to 0.547, demonstrating enhanced long-term prediction accuracy. This extended training, conducted with consistent hyperparameters (learning rate $10^{-3}$, clip range 0.2), strengthened the agent’s ability to anticipate and adapt to dynamic industrial conditions, as evidenced by refined control actions in operational inference scenarios.
The PPO implementation employs a batch size of 64 samples for stable gradient estimation, a learning rate of $10^{-3}$ for efficient policy updates, 10 training epochs per policy update, a discount factor $\gamma=0.99$ for balancing immediate and future rewards, a clipping parameter $\epsilon=0.2$ to limit update magnitude, and a GAE parameter $\lambda=0.95$ for advantage estimation. These settings collectively ensure that policy learning proceeds steadily without catastrophic forgetting or excessive oscillation in performance.
III-D2 Adaptive Exploration Strategy
ANSR-DT implements an adaptive context-aware exploration approach that dynamically adjusts behavior based on operational context, as formalized in Eq. 3. When the system detects deviations through sensor analysis and symbolic reasoning, it switches to exploratory mode by setting the deterministic flag to false, encouraging discovery of new optimal actions. During normal operation, the system defaults to exploitative mode, prioritizing actions with known high rewards [20].
$\pi(a|s)=\begin{cases}\pi_{\text{explore}}(a|s)&\text{while deviation}\\ \pi_{\text{exploit}}(a|s)&\text{otherwise}\end{cases}$ (3)
This context-sensitive strategy ensures that the system balances learning new responses during anomalous conditions with reliable execution during normal operations, avoiding both excessive exploration that could disrupt stable processes and insufficient exploration that would prevent adaptation to novel situations.
III-D3 Dynamic Rule Updating
The Adaptation Layer continuously refines the symbolic knowledge base by incorporating new rules derived from reinforcement learning outcomes. This dynamic updating process ensures that the symbolic reasoning component evolves alongside changing operational conditions and user preferences, maintaining alignment between the system’s explicit knowledge representation and its learned policies. The integration of RL-derived insights into symbolic rules creates a feedback loop where improved policies inform better rules, which in turn guide more effective policy learning.
Figure 2: System operation sequence of the ANSR-DT framework. The diagram illustrates the interaction between components such as the Sensor Network, Data Manager, ML Module, Rule Engine, Digital Twin, and User Interface for real-time time-series pattern extraction and adaptive operations.
III-E Framework Integration and Adaptive Operation
ANSR-DT implements tight integration across its three layers to enable seamless information flow and adaptive decision-making. The operational sequence illustrated in Fig. 2 shows how sensor data progresses through preprocessing, neural pattern extraction, symbolic reasoning, and adaptive control generation in a continuous cycle. This integration maintains bidirectional communication between physical and digital components through a data preprocessing pipeline and feedback control loop.
The data preprocessing pipeline systematically validates and normalizes sensor data before feeding into the processing layer. Sensor readings undergo standardization using rolling statistics to ensure consistent scaling across different measurement types. Multi-sensor data streams are temporally aligned to account for varying sampling rates across temperature, vibration, and pressure sensors. The pipeline identifies deviations in raw sensor readings that may indicate measurement faults or genuine operational anomalies requiring investigation. This preprocessing ensures that the neural component receives clean, aligned inputs suitable for temporal pattern extraction.
The feedback control loop provides real-time optimization of the sensor network itself. The processing layer dynamically adjusts sampling rates based on current system state, increasing measurement frequency when potential anomalies are detected and reducing it during stable operation to conserve resources. Automated sensor calibration occurs based on performance metrics derived from prediction accuracy and consistency checks across redundant measurements. The system continuously performs fault detection and sensor reliability assessment, flagging degraded sensors for maintenance before they compromise monitoring effectiveness.
Continuous adaptation operates through an integrated feedback loop encompassing state assessment and adaptive response. The system continuously monitors performance metrics including accuracy, latency, and resource utilization. Efficiency indices track how effectively the system uses computational and sensing resources. Rule activation patterns and effectiveness are analyzed to identify which symbolic rules provide the most value in different operational contexts. Policy performance is evaluated across operational modes to ensure that learned behaviors generalize appropriately. Based on this assessment, the system coordinates adaptive responses including policy refinement through PPO-based learning, knowledge base evolution via dynamic rule extraction, and sensor network configuration optimization. This integration ensures that ANSR-DT maintains optimal performance while adapting to changing operational conditions, with each layer contributing to the system’s overall adaptability and robustness [28].
III-F Operational Example: Key Event Detection
An operational scenario illustrates ANSR-DT’s integrated decision-making process. The framework detected a pressure reading of 18.0 kPa, which activated neural_rule_1 and neural_rule_4 through the symbolic reasoning layer. The system identified this measurement as a potential operational deviation with 99.81% confidence based on the combined evidence from multiple sensor streams and activated symbolic rules. Extended PPO training allowed the system to refine its control response from conservative adjustments (small temperature modifications, minimal vibration changes) to more decisive actions (larger parameter adjustments aimed at rapidly restoring nominal operating conditions), demonstrating the value of reinforcement learning in industrial control scenarios. This example shows how ANSR-DT integrates neural time-series pattern extraction, symbolic reasoning, and reinforcement learning to detect operational deviations, make interpretable decisions, and adapt control strategies to changing conditions in real-time.
IV Implementation
This section details the technical realization of ANSR-DT, emphasizing methodologies for synthetic data generation, component development, and system integration. The complete implementation, including documentation and example scripts, is available as open-source at https://github.com/sbhakim/ansr-dt.
IV-A Synthetic Data Generation and Rationale
To evaluate ANSR-DT in controlled yet realistic conditions, we developed a synthetic data generation module that simulates industrial manufacturing scenarios. This approach offers perfect ground truth labeling, systematic testing capabilities, and reproducibility for comparative evaluation. The synthetic environment allows validation of system performance without the privacy and security concerns associated with sensitive industrial data, while enabling precise control over the types and frequencies of operational deviations injected for testing.
The dataset comprises 5,000 samples collected at 5-minute intervals, featuring seven variables that capture key aspects of industrial operation. Temperature, vibration, and pressure measurements represent direct sensor readings from physical equipment. Operational hours track system uptime to capture aging and wear effects. The efficiency index quantifies normalized deviation from baseline performance. System state encodes categorical operational modes (startup, normal operation, maintenance, shutdown). Performance score represents inverse efficiency as a derived metric for optimization objectives. Key events, injected at a 5% rate, mimic operational thresholds such as pressure spikes, maintenance events including vibration reductions following component replacement, and state transitions characterized by pressure shifts during mode changes. Events are generated via a multivariate normal distribution with correlations ranging from 0.3 to 0.5 to simulate realistic dependencies between sensor streams. Savitzky-Golay filters [24] smooth the generated signals to remove unrealistic high-frequency noise while preserving the characteristic timescales of industrial processes. This design balances realism with computational efficiency, prioritizing rapid prototyping while features like operational hours and system state enhance contextual modeling for RL and symbolic reasoning.
For validation, we conducted statistical tests including Kolmogorov-Smirnov and Anderson-Darling tests [29], supplemented by cross-correlation analysis to ensure data fidelity. The visual representations in Fig. 3 illustrate the characteristics of our synthetic data. Panel (a) shows vibration measurements over operational hours with labeled event types. Panel (b) displays the distribution of vibration sensor values, distinguishing between normal operational data and key events through color coding. Panel (c) demonstrates pressure readings during a specific time window, highlighting dynamic changes in system state. These visualizations confirm that the synthetic dataset captures realistic patterns and event characteristics suitable for evaluating ANSR-DT’s detection and adaptation capabilities.
Figure 3: Multi-faceted visualization of synthetic sensor data: (a) Vibration measurements (mm/s) over operational hours showing operational thresholds, gradual events, and other key events; (b) Distribution of vibration sensor values distinguishing between normal operational data and key events; (c) Pressure readings (kPa) illustrating dynamic changes in system state during specific time steps. These visualizations demonstrate the realistic patterns and event characteristics incorporated in the synthetic dataset.
IV-B Neuro-Symbolic Reasoning Engine Implementation
IV-B1 Deep Learning Component
The CNN-LSTM neural architecture comprises two Conv1D layers with 64 and 128 filters (kernel size 3), followed by batch normalization and max-pooling. Two LSTM layers with 100 and 50 units then process the spatially-extracted features, with an attention mechanism prioritizing critical time steps within the temporal sequences. This configuration totals 189,481 parameters, carefully balanced to provide sufficient capacity for the 5,000-sample dataset without overfitting on the relatively compact training set.
For optimization, we employed Bayesian optimization [30] to tune hyperparameters including learning rates spanning $10^{-4}$ to $10^{-2}$, dropout rates from 0.1 to 0.5, and network architecture configurations. The final configuration used a learning rate of $10^{-3}$ and dropout rate of 0.3 based on validation performance across the hyperparameter search space. Time series cross-validation with a walk-forward approach ensured robust performance on sequential data, where each fold respects temporal ordering to prevent information leakage from future to past. Early stopping with patience of 10 epochs prevented overfitting by terminating training when validation loss ceased improving. The model was trained using the Adam optimizer with batch size 64 for 20 epochs, achieving convergence as shown in Fig. 5.
IV-B2 Symbolic Reasoning Component
The symbolic component translates neural network outputs into interpretable logical rules and facts using a Prolog-compatible format [27]. The knowledge base contains domain-specific predicates including sensor_reading(Sensor, Value, Time) to encode raw measurements, system_state(State, Time) to represent operational modes, and key_event(Type, Confidence, Time) to flag detected anomalies. These predicates enable logical inference over temporal sequences of observations, supporting queries about system state evolution and event causality.
The framework extracted 14 rules from test predictions, with the rule count stabilizing at 14 across PPO training durations of 10,240 and 200,704 timesteps. This consistency, despite varying RL policies, demonstrates the symbolic component’s reliability in translating neural predictions into interpretable rules on the condensed dataset. Neural rule 1 flags operational deviations with confidence 0.999 when all metrics—temperature, vibration, pressure, efficiency index—fall below their respective thresholds. This rule activated during the inference scenario with pressure reading 18.0, demonstrating actionable symbolic reasoning. Neural rule 6 combines gradient-based conditions (rising efficiency index, temperature, and vibration) with threshold-based conditions, achieving confidence 0.865 but zero test set activations, suggesting it captures rare but potentially important patterns.
neural_rule_1 :-
    feature_threshold(efficiency_index, _, low),
    feature_threshold(pressure, _, low),
    feature_threshold(temperature, _, low),
    feature_threshold(vibration, _, low).
    % Confidence: 0.999, Test Set Activations: 1
neural_rule_6 :-
    feature_gradient(efficiency_index, _, high),
    feature_gradient(temperature, _, high),
    feature_gradient(vibration, _, high),
    feature_threshold(efficiency_index, _, low),
    feature_threshold(pressure, _, low),
    feature_threshold(temperature, _, low),
    feature_threshold(vibration, _, low).
    % Confidence: 0.865, Test Set Activations: 0
For each candidate rule, we calculated confidence scores using the formula shown in Eq. 4, where $\text{support}(r)$ is the frequency of rule $r$’s condition being satisfied, $\text{support}(r\Rightarrow c)$ is the frequency of both the condition and conclusion being satisfied, and $\text{TP}/(\text{TP}+\text{FP})$ is the precision of the rule’s predictions. Rules with confidence scores above threshold $\tau=0.85$ were incorporated into the knowledge base. Conflicting rules were resolved through confidence-based prioritization, where higher-confidence rules take precedence, and specificity analysis, where more specific rules (those with more conditions) override more general rules when both apply. Most learned rules achieved confidence scores above 0.9, demonstrating high reliability in rule extraction.
$\text{conf}(r)=\frac{\text{support}(r\Rightarrow c)}{\text{support}(r)}\times\frac{\text{TP}}{\text{TP}+\text{FP}}$ (4)
IV-B3 Neuro-Symbolic Integration Process
The integration between neural and symbolic components operates through a bidirectional interface ensuring seamless information flow [18]. Neural network outputs are converted into symbolic facts using a thresholding mechanism, where predictions exceeding a specified confidence threshold become symbolic facts available for reasoning. Patterns detected by the neural component are analyzed to extract potential rules, which are then evaluated for confidence against the observed data. Rules exceeding the confidence threshold $\tau$ are added to the knowledge base, subject to conflict resolution as described above.
Symbolic insights remained identical across PPO training runs of 10,240 and 200,704 timesteps, consistently including System Stress Triggered and activations of neural_rule_1 and neural_rule_4. This stability, despite a refined RL policy, enhances interpretability by ensuring reliable, repeatable explanations for human operators—a key advantage in safety-critical settings where decision rationales must be consistent and auditable. Table I summarizes the input and output specifications for each major component, illustrating how data flows through the integrated system.
TABLE I: Component Input/Output Specifications
Component Input Output
Sensor Network Physical measurements Normalized sensor readings
CNN-LSTM Multivariate time series [batch, time_steps, features] Feature vectors and operational deviation probabilities
Symbolic Reasoner Symbolic facts derived from neural output Logical inferences and rule activations
PPO Agent Current state (symbolic facts and sensor readings) Action selection with probability distribution
Rule Extractor Neural patterns and symbolic facts Candidate rules with confidence scores
IV-C Reinforcement Learning Implementation
ANSR-DT implements the PPO algorithm [10] for adaptive decision-making in a custom OpenAI Gym-compatible environment. The environment encapsulates sensor readings, system states, and derived metrics as observation space, while providing operational adjustments as action space. Each state combines raw sensor readings (temperature, vibration, pressure), processed features (operational hours, efficiency index), and symbolic facts (activated rules, current system state) to provide a comprehensive view of the system. The action space includes continuous control parameters such as operational settings and threshold adjustments, and discrete decisions such as maintenance scheduling and mode switching.
The multi-objective reward function balances efficiency, user satisfaction, and safety as formalized in Eq. 5. Efficiency measures how well the system maintains performance metrics near target values. Satisfaction captures alignment with operator preferences and historical best practices. Safety penalizes actions that could lead to hazardous conditions or violate operational constraints. The weighting coefficients $\alpha_{1}=0.5$, $\alpha_{2}=0.3$, and $\alpha_{3}=0.2$ reflect the relative importance of each objective, prioritizing efficiency while ensuring that safety and user preferences are adequately considered.
$\small R(s,a)=\alpha_{1}\cdot\text{Efficiency}(s,a)+\alpha_{2}\cdot\text{Satisfaction}(s,a)+\alpha_{3}\cdot\text{Safety}(s,a)$ (5)
The policy network is implemented as a multi-layer perceptron with three hidden layers containing 256, 128, and 64 units respectively, using ReLU activations. This architecture maps states to action probabilities for discrete actions and action means for continuous actions. The training configuration employed a clipping parameter $\epsilon=0.2$ to constrain policy update magnitude, a discount factor $\gamma=0.99$ for balancing immediate and future rewards, a GAE parameter $\lambda=0.95$ for advantage estimation, and a learning rate of $10^{-3}$ with 10 epochs per policy update and batch size 64.
IV-D Dynamic Rule Updating Mechanism
The continuous rule updating mechanism adapts the symbolic knowledge base based on reinforcement learning insights. Rule candidate generation occurs when the neural component detects novel patterns or when RL agent experiences suggest that existing rules may be incomplete or suboptimal. Each candidate rule’s confidence is assessed based on its predictive accuracy (how often it correctly identifies key events) and support in the observed data (how frequently its conditions are satisfied). Rules exceeding the confidence threshold are incorporated into the knowledge base according to Eq. 6, where existing rules are augmented with new facts derived from recent observations and learning outcomes.
$\text{Rule}_{\text{new}}=\text{Update}(\text{Rule}_{\text{existing}},\text{Fact}_{\text{new}})$ (6)
Conflicting rules are resolved through confidence-based prioritization, where higher-confidence rules take precedence when multiple rules apply to the same situation. Specificity analysis ensures that more detailed rules (those with more conditions) override more general rules to capture nuanced operational contexts. A periodic rule pruning process removes rules with decreasing confidence or relevance, identified through tracking of activation frequency and predictive accuracy over sliding time windows. This dynamic updating ensures that the symbolic reasoning component evolves alongside changing operational conditions and learning insights.
IV-E Application Scenarios
To demonstrate ANSR-DT’s practical utility, we implemented three industrial application scenarios derived from our synthetic data. The predictive maintenance scenario monitors equipment health through temperature and vibration sensors to predict maintenance needs before critical failures occur. The neural component identifies vibration signatures indicative of bearing wear, manifested as increased amplitude in specific frequency bands. Symbolic reasoning provides interpretable maintenance recommendations using rules such as high_frequency_vibration(X) $\land$ rising_temperature(X) $\rightarrow$ maintenance_required(X, urgent). The system learns optimal maintenance scheduling based on operational constraints (minimizing downtime) and historical outcomes (balancing preventive maintenance costs against failure costs).
The process optimization scenario focuses on optimizing operational parameters for efficiency. This implementation integrates temperature, pressure, and derived efficiency metrics to monitor process performance continuously. The RL component learns optimal operational settings under varying conditions, discovering that certain temperature-pressure combinations maximize efficiency for different product types or environmental conditions. Symbolic rules explain why specific settings are optimal in given contexts, enabling operators to understand and validate the system’s recommendations. The system adapts to operator preferences while maintaining safety constraints, ensuring that optimization never compromises operational safety margins.
The deviation detection and response scenario demonstrates ANSR-DT’s capabilities in identifying and responding to operational anomalies. This implementation uses multi-sensor fusion to analyze temperature, vibration, and pressure data for complex event patterns that may not be apparent from any single sensor stream. The system provides graduated responses prioritized by event severity and confidence, ranging from flagging low-confidence anomalies for operator review to automatically triggering protective actions for high-confidence critical events. Continuous learning allows adaptation to new deviation patterns, with symbolic reasoning offering operators clear explanations of detected operational deviations and recommended responses [31].
IV-F Integration and System Architecture
We integrated all components into a cohesive pipeline with well-defined interfaces between layers, as illustrated in Fig. 4. The implementation includes a data preprocessing pipeline for sensor data normalization and fusion, handling missing values through interpolation and detecting measurement outliers through statistical process control. The neural component is implemented in TensorFlow 2.8, leveraging GPU acceleration for training and inference. A symbolic reasoning engine using ProbLog 2.2 [27] handles probabilistic logic inference over the extracted rule base. The reinforcement learning module, based on Stable Baselines3 [32], provides robust implementations of PPO with vectorized environments for efficient training. These components connect through a Python-based middleware with standardized interfaces defined through abstract base classes, ensuring modularity and extensibility. Visualization tools monitor system performance in real-time and explain decisions through interactive knowledge graphs and rule activation traces, as shown in Fig. 8.
Figure 4: ANSR-DT framework architecture with three layers: Physical Environment for sensor integration, Processing Layer for neuro-symbolic reasoning, and Adaptation Layer for reinforcement learning and rule updating. Solid arrows denote data flow, dashed arrows indicate feedback loops for dynamic adaptation.
IV-G Future Extensions
Based on our implementation experience, several promising directions for future enhancements emerge. Scaling to larger datasets will test rule management beyond the current 14 rules, requiring development of hierarchical rule structures to organize large rule bases. Diversifying rule extraction to capture more complex patterns, such as temporal sequences of events or conditional dependencies between sensor streams, will leverage extended PPO training for enhanced adaptability. Adapting the framework for deployment on edge devices with constrained computational resources will require model compression and quantization techniques. Integrating natural language processing will enable more intuitive human-machine interaction, allowing operators to query the system in natural language and receive explanations in accessible terms. Evaluating the framework’s applicability in diverse domains beyond manufacturing, such as energy systems or transportation networks, will assess generalization capabilities [33]. These extensions aim to further enhance the adaptability, interpretability, and practical utility of ANSR-DT for complex industrial applications.
V Results and Analysis
We evaluate ANSR-DT against traditional deep learning approaches using a dataset of 5,000 multivariate time series sequences described in Section IV. Our analysis encompasses classification performance, feature importance, rule extraction efficacy, and component contribution analysis to demonstrate the framework’s capabilities in dynamic event detection and adaptability.
V-A Experimental Setup
Experiments were conducted using synthetic industrial data with a 60/20/20 train/validation/test split containing 5% labeled dynamic patterns distributed across operational thresholds, maintenance patterns, state transitions, and interaction deviations. To address class imbalance in the dataset, we applied class weighting (0: 64.7, 1: 0.5) during model training, ensuring that the minority class (key events) received appropriate emphasis during gradient computation. We established a rigorous comparison framework against CNN-LSTM baseline models to isolate the contributions of the neuro-symbolic integration and reinforcement learning components.
V-B Performance Analysis and Time-Series Pattern Extraction
Table II presents comparative performance metrics between ANSR-DT and the CNN-LSTM baseline. Our framework achieves a peak validation accuracy of 99.5% by epoch 15 across 20 training epochs, compared to 86.87% for the CNN-LSTM baseline. However, fluctuations observed during training (e.g., validation accuracy dipping to 55% at epoch 16) suggest potential overfitting on the reduced dataset, mitigated through early stopping and class weighting to address imbalance. The training progression illustrated in Fig. 5 shows convergence with validation accuracy reaching its peak at epoch 15, indicating efficient learning despite some fluctuations attributable to the smaller dataset size.
TABLE II: Performance Comparison: ANSR-DT vs CNN-LSTM Baseline
Metric ANSR-DT CNN-LSTM Improve. (%) p-value
Accuracy (%) 99.5 (peak) 86.87 +12.63 << 0.01
Precision 0.80 $\pm$ 0.05 0.9358 -14.50 << 0.05
Recall 0.75 $\pm$ 0.04 0.6355 +17.82 << 0.05
F1-Score 0.77 $\pm$ 0.03 0.7570 +1.58
ROC-AUC 0.85 $\pm$ 0.03 0.8049 +5.58 << 0.01
Figure 5: Training and validation accuracy and loss trends over 20 epochs, showing convergence with some fluctuations. Validation accuracy reached a peak of 99.5% by epoch 15, with class weighting (0: 64.7, 1: 0.5) addressing imbalance in the dataset.
The most substantial improvement appears in recall performance, with a 17.82% increase ($p<0.05$) over the baseline, indicating significantly enhanced capability to identify critical patterns with fewer false negatives—a crucial characteristic for industrial safety applications where missing operational deviations can lead to equipment damage or safety hazards. This high detection accuracy is achieved through the combination of CNN-LSTM architecture with an attention mechanism, which effectively learns temporal dependencies across multiple sensor streams. The precision-recall trade-off, deliberately engineered to prioritize recall in safety-critical industrial settings, reflects the operational reality that false negatives typically carry higher operational risk than false positives. While precision decreased by 14.50%, this trade-off is acceptable given that false alarms can be reviewed by operators, whereas missed critical events may propagate into serious failures.
Fig. 6 provides visual confirmation of classification performance across decision thresholds. The confusion matrix in panel (a) shows high true positive and true negative rates, with 343 true negatives (100.0%) and 153 true positives (98.1%), minimal false negatives (3 samples, 1.9%), and no false positives. This distribution confirms the model’s ability to maintain high specificity while achieving strong sensitivity. The precision-recall curve in panel (b) demonstrates robust performance across different decision thresholds with an Average Precision score of 1.00, indicating that the model maintains high precision even as recall increases—a desirable characteristic for operational deployment where threshold selection must balance detection sensitivity with alarm rate.
[…]
Figure 6: Performance Metrics: (a) Confusion Matrix showing classification accuracy with 343 true negatives (100.0%) and 153 true positives (98.1%), with minimal false negatives (1.9%) and no false positives; (b) Precision-Recall Curve with an Average Precision (AP) score of 1.00, demonstrating near-perfect performance across decision thresholds.
V-C Feature Importance and Rule Extraction Analysis
Feature importance analysis, depicted in Fig. 7, identifies vibration as the most influential variable in dynamic event detection, with an importance score of 0.000175, followed by temperature (0.000155) and pressure (0.000115). This ranking highlights the critical role of vibration in capturing early signs of system deviations, which may precede changes in temperature or pressure in industrial systems where mechanical degradation manifests first through altered vibration signatures before affecting thermal or hydraulic characteristics. Operational hours and efficiency index also contribute significantly, with scores of 0.000085 and 0.000075 respectively, indicating their relevance in predictive modeling through capturing cumulative wear effects and performance drift. System state and performance score exhibit lower importance, suggesting that while they provide supplementary context, they are less critical for event detection in this particular scenario. These results underscore the value of multivariate analysis in uncovering complex relationships that single-sensor approaches might overlook, validating the CNN-LSTM architecture’s ability to automatically learn feature importance rather than relying on manual feature engineering.
Figure 7: Feature Importance Analysis showing vibration as the top predictor (0.000175), followed by temperature (0.000155) and pressure (0.000115), highlighting the framework’s ability to capture complex multivariate relationships.
The rule extraction mechanism successfully extracted 14 rules from test predictions, with 64% (9 rules) exceeding a 0.9 confidence score. These rules include both threshold-based conditions (e.g., neural_rule_1, which activates when all sensor readings fall below specified thresholds) and gradient-based conditions (e.g., neural_rule_7, which responds to rates of change in sensor readings). The rules remained stable across PPO training durations of 10,240 and 200,704 timesteps, reflecting robust symbolic reasoning on the condensed dataset. This stability across varying RL policies highlights the symbolic component’s reliability in translating neural predictions into interpretable rules. The consistency demonstrates that the core patterns detected by the neural network—and subsequently encoded as symbolic rules—represent fundamental operational relationships rather than artifacts of specific policy configurations.
The knowledge graph generation process, illustrated in Fig. 8, creates a directed graph structure connecting sensor readings, system states, activated rules, and generated insights. The visualization shows how a critical system event is diagnosed through the reasoning chain. The central State node (green) indicates ‘Critical’ status, directly linked to poor performance Metrics (gray; E=0.07, P=7.2) and a high-confidence key event (salmon; Severity 2, Confidence 1.0). Contributing sensor readings (blue; Temperature, Vibration, Pressure) connect via ‘related_to’ edges (dashed gray) to high-impact symbolic rules (pink; e.g., neural_rule_1 with confidence 1.00). These rules ‘detect’ the operational deviation (dashed red edges). Resulting Insights (yellow), such as ‘System Stress’ and rule activations, are ‘explained_by’ the deviation (solid green edges) or ’generated’ by rules (dotted purple edges). This visualization provides operators with a comprehensive view of the decision-making process, highlighting how sensor data leads to specific insights through rule activation paths, thereby enhancing trust and enabling validation of system conclusions.
Figure 8: Focused ANSR-DT Knowledge Graph snapshot illustrating the diagnosis of a critical system event. The central State node (green) indicates “Critical” (C), directly linked to the poor performance Metrics (gray; E=0.07, P=7.2) and the high-confidence key event detected (salmon; S=Severity 2, C=Confidence 1.0). Contributing sensor readings (blue; Temp, Vibr, Pres) are connected via “related_to” edges (dashed gray) to a filtered set of high-impact symbolic rules (pink; e.g., nr_1 C:1.00). These rules “detect” the operational deviation (dashed red edges). Resulting Insights (yellow), such as “System Str” (System Stress) and rule activations (“neural_rule_X”), are shown as being “explained_by” the deviation (solid green edges) or “generated” by rules (dotted purple edges). This focused view demonstrates the framework’s interpretability by visualizing the reasoning chain from sensor inputs, through symbolic rules, to the critical state diagnosis and associated insights. Node colors differentiate types (legend omitted for brevity, explained in text/standardized).
V-D Component Contribution Analysis
Ablation studies underscored the rule-learning component’s critical role in overall system performance. Removing the symbolic reasoning component caused a 7.4% F1-score drop, highlighting its contribution to both time-series pattern extraction and interpretability. This degradation occurs because the symbolic component provides structured feedback to the neural network through rule-based constraints, guiding attention toward operationally relevant patterns rather than spurious correlations. The combination of attention mechanism and rule learning demonstrated the strongest synergistic effect, particularly in operational deviation detection accuracy. However, some improvements were not statistically significant at higher noise levels (signal-to-noise ratios exceeding 30dB), identifying a potential area for robustness enhancement through advanced preprocessing or more sophisticated attention mechanisms that can filter noise more effectively.
V-E Adaptability Assessment
ANSR-DT demonstrated superior adaptability with extended PPO training increasing the explained variance from 0.447 to 0.547, representing a 22.4% improvement in the agent’s ability to predict value functions accurately. This improvement in adaptation capability stems from the integrated neuro-symbolic approach and continuous rule updating mechanism working in concert. The framework successfully identified and adapted to the injected operational deviations, with particularly strong performance on operational threshold violations (98.2% detection rate) and interaction pattern deviations (95.7% detection rate), though performance on gradual drift patterns remained lower (87.3% detection rate), suggesting an area for future improvement.
The neuro-symbolic integration allowed the framework to update rules based on observed patterns and reinforcement learning outcomes in a closed-loop fashion. When encountering new deviation patterns during training, the system required relatively few iterations (typically 5-10 policy updates) to develop high-confidence rules, demonstrating efficient learning compared to pure neural approaches that often require substantially more training samples. The symbolic representation facilitated faster adaptation than pure neural approaches by enabling targeted rule modifications rather than full model retraining—when a new pattern emerges, the system can add or modify specific rules rather than recalibrating the entire neural network weight matrix.
Extended PPO training to 200,704 timesteps adjusted the recommended action for operational deviations from more conservative values (temperature adjustment: -2.3, vibration adjustment: +0.8, pressure adjustment: -1.1) to bolder responses (temperature adjustment: -4.7, vibration adjustment: +1.9, pressure adjustment: -2.8), indicating policy refinement toward more aggressive corrective actions. This shift reflects RL’s ability to refine control strategies over time while maintaining high deviation detection confidence (99.81%), demonstrating that the learned policy increasingly trusts its ability to make larger adjustments without destabilizing the system.
V-F Limitations and Future Research Directions
Despite ANSR-DT’s strong performance, several limitations suggest directions for future research. The reduced rule count (14 rules) in this study limits scalability testing of the rule management system, which previously faced challenges beyond 50 rules in preliminary experiments. Future exploration with larger datasets containing more diverse operational patterns will be essential to test whether the rule extraction and conflict resolution mechanisms scale effectively to complex industrial environments with hundreds of potential rules. Neuro-symbolic integration adds computational overhead, requiring approximately 35% more processing time compared to pure neural approaches, necessitating optimization for real-time use in industrial applications. This represents a trade-off for the interpretability benefits—the additional computation enables transparent decision-making but may challenge deployment in latency-sensitive applications.
Promising research directions include development of hierarchical rule structures for improved scalability, where high-level rules capture general operational principles and lower-level rules encode specific context-dependent responses. Optimization of the rule validation process through incremental validation, where only modified rules are re-validated rather than the entire rule base, and caching techniques that store frequently-used inference results will reduce computational overhead. Enhanced noise resilience through advanced preprocessing methods such as adaptive filtering and more robust feature extraction techniques that explicitly model sensor measurement uncertainty will improve performance in high-noise environments. Future work will also explore expanding beyond the current sensor configuration to incorporate additional sensor modalities including acoustic emissions, electrical signatures, and process flow measurements to capture more complete operational state representations.
VI Discussion
ANSR-DT demonstrates significant advantages in enhancing interpretability and decision-making accuracy through the integration of symbolic reasoning with digital twin technology. The neural components achieve high accuracy (up to 99.5%) in dynamic event detection, validating the CNN-LSTM architecture’s effectiveness across multiple deviation types including operational thresholds, maintenance events, and state transitions. The framework’s ability to extract and maintain 14 rules with consistent confidence scores represents a significant advancement in translating neural insights into human-interpretable knowledge. This capability directly addresses the black-box problem that has limited adoption of AI-enhanced digital twins in safety-critical industrial environments [22], where operators require transparent explanations to validate system recommendations and maintain appropriate situational awareness.
The improved explained variance (0.547) with extended PPO training demonstrates ANSR-DT’s capability to respond to changing conditions without requiring comprehensive retraining of the entire system. Symbolic insights remained identical across PPO training runs of 10,240 and 200,704 timesteps, consistently including System Stress Triggered and activations of neural_rule_1 and neural_rule_4. This stability, despite refined RL policies that learned to take increasingly decisive control actions, enhances interpretability by ensuring reliable, repeatable explanations for human operators—a key advantage in safety-critical settings where decision rationales must be consistent and auditable across different operational phases and policy refinement iterations.
Despite these achievements, several key challenges remain that limit the framework’s full potential in large-scale industrial deployment. Rule extraction efficiency requires substantial improvement to scale beyond the current 14 rules extracted from the condensed dataset. Preliminary experiments with larger rule sets (beyond 50 rules) encountered performance degradation due to increased inference time and potential rule conflicts that the current conflict resolution mechanism struggles to handle efficiently. The current knowledge graph structure constrains representation of complex relationships in industrial systems with numerous interdependent components, suggesting that hierarchical or modular knowledge representations may be necessary for systems with hundreds of interacting subsystems. The neural-symbolic integration introduces computational overhead (approximately 35% increase in processing time), creating latency challenges for real-time processing in resource-constrained environments [9]. This overhead stems primarily from the bidirectional communication between neural and symbolic components, where neural outputs must be converted to symbolic facts and symbolic constraints must be propagated back to guide neural attention.
Environmental noise sensitivity presents an additional limitation, particularly in sensor-dense scenarios where signal quality may vary due to electromagnetic interference, mechanical vibrations, or environmental conditions. Our ablation studies indicate that this sensitivity is most pronounced when the attention mechanism is removed, suggesting that the current attention implementation provides crucial filtering capabilities that could be further enhanced in future iterations through explicit noise modeling or adversarial training approaches that expose the system to realistic noise profiles during development. The framework’s current performance on gradual drift patterns (87.3% detection rate) lags behind its performance on abrupt threshold violations (98.2% detection rate), indicating that the temporal modeling may benefit from longer time windows or explicit drift detection mechanisms.
These limitations primarily stem from the fundamental challenge of bridging connectionist and symbolic paradigms while maintaining both learning capacity and interpretability. While ANSR-DT establishes a foundation for adaptive decision-making in digital twins, addressing these challenges through hierarchical knowledge representations, optimized inference algorithms, and enhanced noise resilience mechanisms is crucial for realizing the full potential of neuro-symbolic integration in complex industrial applications with diverse operational patterns and stringent real-time requirements.
VII Conclusion
ANSR-DT achieves 99.5% peak validation accuracy and extracts 14 stable symbolic rules, establishing a robust foundation for interpretable digital twins in industrial applications. The framework integrates CNN-LSTM neural networks for temporal pattern extraction, Prolog-based symbolic reasoning for transparent decision processes, and PPO reinforcement learning for continuous adaptation. Extended PPO training to 200,704 timesteps improved explained variance from 0.447 to 0.547, demonstrating enhanced long-term prediction accuracy and policy refinement. The symbolic component remained stable across varying training durations, consistently identifying operational deviations and providing reliable explanations for human operators. This combination of high performance and interpretability addresses the black-box problem that has limited AI adoption in safety-critical industrial environments where transparent decision-making is essential.
Future work will focus on three key directions to enhance scalability and robustness. First, developing hierarchical rule structures and optimized inference algorithms will enable management of larger rule bases beyond the current 14 rules, supporting more complex industrial scenarios. Second, enhanced noise resilience through advanced preprocessing and robust feature extraction will improve performance in high-noise industrial environments. Third, expanding sensor modalities and evaluating the framework across diverse domains beyond manufacturing will assess generalization capabilities. These advancements will further bridge the gap between adaptive intelligence and interpretable decision-making, advancing the development of trustworthy industrial digital twin technology.